### v1

# MSUHEP-24-019

Abstract
The Hessian method is widely used in the global analysis of parton distribution functions (PDFs). It uses a set of orthogonal eigenvectors to calculate the first-order derivative terms of a physics observable, so that the leading order PDF-induced uncertainty can be estimated accordingly. In this article, we report an improved Hessian method which introduces the second-order derivatives in the uncertainty estimation. To estimate the effect of the second-order uncertainties, a pseudo global analysis is designed. We also find that the PDF-induced uncertainties on physics variables can be significantly different by taking the second-order effects into account, particularly when the precisions of experimental inputs used in the PDF global analysis are limited.
I Introduction
In the past decades, a global analysis method based on quantum chromo-dynamics (QCD) has been established to study the proton structure. It sufficiently analyzes comprehensive data results from the deep inelastic scattering experiments with the energy scale of a few GeV, to the processes at hadron colliders at $\mathcal{O}(100)$ GeV. In practice, the global analysis is required to not only give the central predictions on parton densities corresponding to the best fit of the data, but also provide a way that the uncertainties can be easily calculated and extrapolated for any particular physics observable. The Hessian method [1] is one of the most widely used strategies to estimate the PDF uncertainties, such as in the latest CT18NNLO and MSHT20 PDFs [2, 3] which use both the DIS and hadron collider data. In this approach, a group of error PDF sets are provided, which represent the possible variations on the parameters of the non-perturbative functions of the quarks and gluons. These error sets correspond to orthogonal eigenvectors in the original parameter space, so that the uncertainty of a particular physics observable can be calculated using the error sets without complicated analysis on the correlations between the original parameters.
However, in the current Hessian method, only the leading order terms in the expansion with respect to a physics observable of those eigenvectors are taken into account to estimate the PDF-induced errors. As more data inputs are included in the global analysis, and more parameters are used in the non-perturbative parameterization of the PDFs, the higher order terms in the expansion of a physics observable may become more important. This has been mentioned in a few previous studies. For example, in the Error PDF Updating Method Package [4], contributions from the second-order terms in the expansion of a single eigenvector, appearing as the diagonal terms in the Hessian matrix as to be discussed later, were estimated based on the CT14 PDFs [5]. Based on the same PDFs, it was found later in Ref. [6] that the higher order terms of a single eigenvector can be sizable in the analysis of the CMS di-jet measurement. However, the off-diagonal terms corresponding to the expansion of multiple eigenvectors cannot be faithfully tested in the above studies since CT14 did not provide the needed error sets for this kind of tasks.
In this work, we introduce an improved Hessian method, which generated the PDF error sets allowing a complete second-order calculation of the uncertainty of a physics observable. To make a detailed comparison between the improved method and the original one, a pseudo global fit is studied. As will be demonstrated in this test, when including the second-order terms, the uncertainty on parton densities could be noticeably different from the original leading order estimation. Such difference can be reduced as the precision of the input data in the global analysis improves. However, the required precision is beyond many experimental inputs included in the present PDF global analysis, especially from those old DIS and fixed-target Drell-Yan data [2]. Hence, it would be important to consider the higher order terms in the uncertainty estimation of the PDF global analysis.
II Method
The Hessian method [1] is designed based on the following expansion:
$\mathcal{O}(z_{i})=\mathcal{O}(z^{0}_{i})+\sum_{i}\frac{\partial\mathcal{O}}{\partial z_{i}}\Big{|}_{z^{0}_{i}}\text{d}z_{i}+\frac{1}{2}\sum_{i,j}\frac{\partial^{2}\mathcal{O}}{\partial z_{i}\partial z_{j}}\Big{|}_{z^{0}_{i},z^{0}_{j}}\text{d}z_{i}\text{d}z_{j}+\cdot\cdot\cdot$ (1)
where $\mathcal{O}$ is a physical observable whose theory prediction depends on PDFs. $\{z_{i}\}$ represent the parameters in the PDFs, and $\{z^{0}_{i}\}$ are the corresponding central values determined as the best fit to the data inputs. When $z_{i}$ is independent of $z_{j}$, for $i\neq j$, the leading order uncertainty on $\mathcal{O}$ can simply be calculated as:
$\delta\mathcal{O}\approx\sqrt{\sum_{i}\left(\frac{\mathcal{O}(z^{0}_{i}+\delta z_{i})-\mathcal{O}(z^{0}_{i}-\delta z_{i})}{2}\right)^{2}}$ (2)
where $\delta z_{i}$ is the uncertainty on $z_{i}$. $\mathcal{O}(z^{0}_{i}\pm\delta z_{i})$ are the calculations of $\mathcal{O}$ where the value of the $i$-th parameter is shifted by ±$\pm\delta z_{i}$, while all other parameters $z_{j}(j\neq i)$ are fixed at their central values $z^{0}_{j}$. To enable the calculation, the global analysis provides a central PDF set $S_{0}=S(z^{0}_{i})$ which can be used to calculate the central prediction of $\mathcal{O}(S_{0})$, and a group of error PDF sets $S_{i\pm}=S(z^{0}_{i}\pm\delta z_{i};z_{j}=z^{0}_{j})$ for $j\neq i$. In this strategy, the first-order derivatives are represented by
$\frac{\partial\mathcal{O}}{\partial z_{i}}\Big{|}_{z^{0}_{i}}=\frac{\mathcal{O}(z^{0}_{i}+\delta z_{i})-\mathcal{O}(z^{0}_{i}-\delta z_{i})}{2\delta z_{i}},$ (3)
while the higher order derivatives $\partial^{2}\mathcal{O}/\partial z_{i}\partial z_{j}$ are ignored. As discussed above, Ref. [4] and Ref. [6] studied effects on $\partial^{2}\mathcal{O}/\partial z^{2}_{i}$.
We first review how to obtain the orthogonal eigenvectors $\{v_{i}\}$, with $\{z_{i}\}$ being the displacement (in the space of $\{z_{i}\}$) along the direction of $\{v_{i}\}$, by the Hessian method. The non-perturbative functions which describe the parton densities at the scale of $Q_{0}$, taken from the CT18NNLO PDF as an example, are written in the formalism as:
$f_{q}(x,Q_{0})=a_{0}x^{a_{1}-1}(1-x)^{a_{2}}P_{q}(x;a_{3},a_{4},\cdot\cdot\cdot)$ (4)
where $x$ is the momentum fraction of the parton $q$ inside the proton. $P_{q}$ is a polynomial designed to allow the parton density to change its shape in order to match the data inputs in different $x$ regions. $\{a_{i}\}$ are the original parameters being determined in the global fit, by minimizing the $\chi^{2}$ defined as:
$\chi^{2}=\sum_{\alpha}\frac{[D_{\alpha}-T_{\alpha}(a_{i})]^{2}}{\sigma^{2}_{\alpha}}$ (5)
where $D_{\alpha}$ and $\sigma_{\alpha}$ are the experimental inputs and their corresponding uncertainties. $T_{\alpha}(a_{i})$ is the theory prediction of $D_{\alpha}$, depending on the PDF shape parameters $\{a_{i}\}$. In practice, $\chi^{2}_{\text{total}}$ is more complicated, which may include correlated systematic uncertainties. Furthermore, to deal with the potential biases of the theory calculation and the experimental measurements, possible inconsistency between data, and the uncertainty due to the choice of the non-perturbative parameterizations, some dynamical and/or global tolerances are also introduced to define the errors of the fitted PDF parameters, such as in the CT18 and MSHT20 [2, 3]. For simplicity, we ignore all these variants in this work as they will not strongly alter our conclusions. In this work, we treat the pseudo-PDF uncertainty to be uncorrelated and purely statistical.
Given the fact that the best fit values of $\{a_{i}\}$ appear near the minimal $\chi^{2}$, we have:
$\chi^{2}=\chi^{2}_{\text{min}}+\frac{1}{2}\sum_{i,j}\frac{\partial^{2}\chi^{2}}{\partial a_{i}\partial a_{j}}(a_{i}-a^{0}_{i})(a_{j}-a^{0}_{j})+\cdot\cdot\cdot.$ (6)
The Hessian matrix is then defined as $H_{ij}=\partial^{2}\chi^{2}/\partial a_{i}\partial a_{j}$. Note that the original parameters $a_{i}$ can be correlated in the global fit because experimental observables usually relate to multiple PDF shape parameters. One can convert $\{a_{i}\}$ to another set of orthogonal parameters $\{z_{i}\}$ by diagonalizing $H_{ij}$:
$\sum_{j}H_{ij}v_{jk}$ $=$ $\epsilon_{k}v_{ik},$
where $\epsilon_{k}$ are the eigenvalues and $v_{ik}$ are the orthogonal and complete eigenvectors. Then, the displacements of $\{a_{i}\}$ are given as
$a_{i}-a^{0}_{i}$ $=$ $\sum_{k}v_{ik}s_{k}z_{k},$ (7)
where the scaling factor $s_{i}$ is introduced to normalize $z_{i}$. Usually, $s_{k}$ is chosen to make
$\Delta\chi^{2}=\chi^{2}-\chi^{2}_{\text{min}}=\sum_{i}z^{2}_{i},$ (8)
so that the best fit of $\chi^{2}$ (denoted as $\chi^{2}_{\text{min}}$) corresponds to $\{z_{i}=0\}$, and the one-standard deviation uncertainty in the direction of the $i$-th eigenvector corresponds to $\{z_{i}=\pm 1;z_{j}=0$ for $j\neq i\}$. By this requirement, $s_{i}$ are approximately equal to $\sqrt{2/\epsilon_{i}}$.
The central PDF set and the error sets can be established. The central PDF set corresponds to the best fit parameters $S_{0}=S(a_{i}=a^{0}_{i})$, or $S_{0}=S(z_{i}=0)$. For the error PDF sets, we denote the ±$\pm 1$ standard variation in the direction of the $i$-th eigenvector as
$S_{i\pm}=S(z_{i}=\pm 1;z_{j}=0(j\neq i)).$ (9)
Hence, the derivative in Eq. (3) of any physical observable $\mathcal{O}$ can be calculated using the PDF error sets:
$\frac{\partial\mathcal{O}}{\partial z_{i}}\Big{|}_{0}=\frac{\mathcal{O}(S_{i+})-\mathcal{O}(S_{i-})}{2},$ (10)
so that the leading order uncertainty is acquired. $S_{0}$ and $S_{i\pm}$ are the PDF sets provided by the global analysis groups such as CTEQ-TEA and MSHT. They can be used to calculate the first-order derivatives in Eq. (1), but not those off-diagonal terms.
Below, we show how to obtain the PDF error sets for calculating the full second-order derivatives in Eq. (1). Following the definition of the derivative, the needed error sets are:
$S^{i\pm}_{j\pm}=S(z_{i}=\pm 1,z_{j}=\pm 1;z_{k}=0\ \mathrm{for}\ k\neq i,j),$ (11)
in which two displacements are varied simultaneously, so that the second-order derivatives in Eq. (3) can be calculated as:
$\frac{\partial^{2}\mathcal{O}}{\partial z_{i}\partial z_{j}}$ $=$
$\frac{1}{2}$$\times$ $\left(\frac{\mathcal{O}(S^{i+}_{j+})-\mathcal{O}(S^{i-}_{j+})}{2}-\frac{\mathcal{O}(S^{i+}_{j-})-\mathcal{O}(S^{i-}_{j-})}{2}\right)$ (12)
Note that for $i=j$, the second-order error sets are $S_{i++}=S(z_{i}=2;z_{k}=0$ for $k\neq i)$ and $S_{i--}=S(z_{i}=-2;z_{k}=0$ for $k\neq i)$. Consequently, we have
$\frac{\partial^{2}\mathcal{O}}{\partial z^{2}_{i}}$ $=$
$\frac{1}{2}$$\times$ $\left(\frac{\mathcal{O}(S_{i++})-\mathcal{O}(S_{0})}{2}-\frac{\mathcal{O}(S_{0})-\mathcal{O}(S_{i--})}{2}\right)$ (13)
According to Eq. (II), the values of the original parameters $\{a_{i}\}$ corresponding to $\{z_{i}=\pm 1,z_{j}=\pm 1;z_{k}=0$ for $k\neq i,j\}$ can be calculated, and thus the second-order error sets $S^{i\pm}_{j\pm}$ can be directly generated. Since the Hessian matrix comes from the derivatives of $\chi^{2}$, the second-order error sets can only be acquired together with the first-order error sets in a complete PDF global analysis.
With full second-order derivatives included, the uncertainty on $\mathcal{O}$ is written as:
$\left(\delta\mathcal{O}\right)^{2}=\sum_{i}\left(\frac{\partial\mathcal{O}}{\partial z_{i}}\right)^{2}+\frac{1}{2}\sum_{i}\left(\frac{\partial^{2}\mathcal{O}}{\partial z^{2}_{i}}\right)^{2}+\frac{1}{2}\sum_{i\neq j}\left(\frac{\partial^{2}\mathcal{O}}{\partial z_{i}\partial z_{j}}\right)^{2},$ (14)
where the first, second and third terms correspond to the contributions from the first-order derivatives, the second-order diagonal derivatives, and the second-order off-diagonal derivatives in Eq. (1). The process of proof of Eq. (14) is given in the Appendix. One can also calculate the uncertainty in an alternative way:
i) give each displacement $z_{i}$ (along each eigenvector $v_{i}$) a random value according to the normalized gaussian distribution $G(0,1)$.
ii) calculate $\mathcal{O}=\mathcal{O}(S_{0})+\sum_{i}\frac{\partial\mathcal{O}}{\partial z_{i}}z_{i}+\frac{1}{2}\sum_{i,j}\frac{\partial^{2}\mathcal{O}}{\partial z_{i}\partial z_{j}}z_{i}z_{j}$, where the derivatives are calculated using the error sets.
iii) repeat step i) and ii) many times, and interpret the many resulting values of $\mathcal{O}$ as a statistical distribution of $\mathcal{O}$, with its one standard deviation representing the PDF-induced uncertainty of the observable $\mathcal{O}$.
Since the first- and second-order derivatives (i.e. $\frac{\partial\mathcal{O}}{\partial z_{i}}$ and $\frac{\partial^{2}\mathcal{O}}{\partial z_{i}\partial z_{j}}$) can be readily calculated, the above bootstrap procedure can be quickly done.
The bootstrap procedure gives consistent uncertainty with Eq. (14). It is also convenient to calculate the correlation of the PDF-induced uncertainties between two physical observables $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$. One can calculate $(\delta\mathcal{O}_{1})^{2}$, $(\delta\mathcal{O}_{2})^{2}$ and $(\delta(\mathcal{O}_{1}+\mathcal{O}_{2}))^{2}$ via either aforementioned method. The correlation is then easily acquired on the basis of its original definition of correlation:
$\rho(\mathcal{O}_{1},\mathcal{O}_{2})=\frac{(\delta(\mathcal{O}_{1}+\mathcal{O}_{2}))^{2}-(\delta\mathcal{O}_{1})^{2}-(\delta\mathcal{O}_{2})^{2}}{2\delta\mathcal{O}_{1}\delta\mathcal{O}_{2}}$ (15)
III Pseudo global analysis
In this section, we introduce a pseudo global analysis to numerically test the improved Hessian method. We start with a set of parton densities taken from the CT18NNLO PDFs [2]. For valence $u_{V}$ and $d_{V}$, the non-perturbative functions are:
$f_{q}(x)$ $=$ $a_{0}x^{a_{1}-1}(1-x)^{a_{2}}P^{V}_{a}(y)$
$P^{V}_{a}(y)$ $=$ $\sinh[a_{3}](1-y)^{4}+\sinh[a_{4}]4y(1-y)^{3}+\sinh[a_{5}]6y^{2}(1-y)^{2}$ (16)
$+\left(1+\frac{1}{2}a_{1}\right)4y^{3}(1-y)+y^{4},$
where $y\equiv\sqrt{x}$. For both $u_{V}$ and $d_{V}$, $a_{0}$ is fixed, following the flavor sum rules. We further require $a^{u_{V}}_{1}=a^{d_{V}}_{1}$ and $a^{u_{V}}_{2}=a^{d_{V}}_{2}$, as done in CT18 global analysis. For the sea quarks $\bar{u}$, $\bar{d}$, and $s=\bar{s}$, the non-perturbative functions are:
$f_{q}(x)$ $=$ $a_{0}x^{a_{1}-1}(1-x)^{a_{2}}P^{\text{sea}}_{a}(y)$
$P^{\text{sea}}_{a}(y)$ $=$ $(1-y)^{5}+a_{4}5y(1-y)^{4}+a_{5}10y^{2}(1-y)^{3}$ (17)
$+a_{6}10y^{3}(1-y)^{2}+a_{7}5y^{4}(1-y)+a_{8}y^{5},$
where $y\equiv 1-(1-\sqrt{x})^{a_{3}}$. We require $a^{\bar{u}}_{0}=a^{\bar{d}}_{0}$, $a^{\bar{u}}_{1}=a^{\bar{d}}_{1}=a^{s}_{1}$, $a^{\bar{u}}_{2}=a^{\bar{d}}_{2}$, $a^{\bar{u}}_{3}=a^{\bar{d}}_{3}=a^{s}_{3}=4$, $a^{\bar{d}}_{8}=a^{s}_{8}=1$, $a^{s}_{4}=a^{s}_{5}$, and $a^{s}_{6}=a^{s}_{7}$, following CT18. The distribution of the gluon is ignored in this test, since it would complicate the PDF analysis to add gluon contributions in the pseudo data. The charm quark is also ignored because it is highly correlated with the gluon. Therefore, we have 24 free parameters in total. Their central values are also set to the best fit of CT18NNLO at the initial energy scale $Q_{0}=1.3$ GeV [2], as listed in Tab. 1. Figure 1 depicts $xf_{q}(x)$ for different quarks. We note that in this study, these pseudo PDFs will not be evolved to a high energy scale relevant to the pseudo data introduced below. These pseudo PDFs do not resemble any of the CT18 PDF sets. They are merely designed to simplify our discussions on the effect of the second-order derivatives in Eq. (1) on the estimates of PDF-induced uncertainties.
$u_{V}$ $d_{V}$ $\bar{u}$ $\bar{d}$ $s=\bar{s}$
$a_{0}$ 3.385 (fixed) 0.490 (fixed) 0.414 0.414 0.288
$a_{1}$ 0.763 0.763 -0.022 -0.022 -0.022
$a_{2}$ 3.036 3.036 7.737 7.737 10.31
$a_{3}$ 1.502 2.615 4 (fixed) 4 (fixed) 4(fixed)
$a_{4}$0.147 1.828 0.618 0.292 0.466
$a_{5}$ 1.671 2.721 0.195 0.647 0.466
$a_{6}$ 0.871 0.474 0.225
$a_{7}$ 0.267 0.741 0.225
$a_{8}$ 0.733 1 (fixed) 1 (fixed)
Table 1: Values of parameters in the non-perturbative functions, taken from CT18NNLO.
Figure 1: $xf_{q}(x)$ for $u_{V}$, $d_{V}$, $\bar{u}$, $\bar{d}$ and $s=\bar{s}$ in the pseudo global analysis.
A group of pseudo observables are also constructed according to the PDFs in Eq. (III) and Eq. (III). These observables, with uncertainties assigned, will be used as the data input in the pseudo global analysis. First, we introduce $\sigma^{\text{DIS}}_{\gamma}(x)$ and $\sigma^{\text{DIS}}_{Z}(x)$:
$\sigma^{\text{DIS}}_{\gamma}(x)$ $=$ $u_{V}(x)+2\bar{u}(x)+0.25\left[d_{V}(x)+2\bar{d}(x)+2s(x)\right]$
$\sigma^{\text{DIS}}_{Z}(x)$ $=$ $u_{V}(x)+2\bar{u}(x)+1.2\left[d_{V}(x)+2\bar{d}(x)+2s(x)\right]$ (18)
as pseudo observables similar to the cross section measurements of the photon-exchange DIS process $\ell+p\rightarrow\gamma^{*}\rightarrow\ell+X$, and the $Z$-exchange DIS process $\ell+p\rightarrow Z/\gamma^{*}\rightarrow\ell+X$. The $q(x)$ functions are taken from Eq. (III) and Eq. (III), with $\{a_{i}\}$ values from Tab. 1. The coefficient 0.25 (1.2) roughly reflects the strength difference between the down-type and up-type quarks coupling to $\gamma$ ($Z$). Note that we take the definitions of $u\equiv u_{V}+\bar{u}$ and $d\equiv d_{V}+\bar{d}$, so that the $\bar{u}$ and $\bar{d}$ terms have a coefficient of 2. For simplicity, we ignore other contributions in a real calculation of the DIS process, such as the evolution of PDFs as a function of $Q$, the electroweak couplings and the higher order corrections. Although these effects would significantly change the predictions on any physical observable, the main sensitivity to the particular parton information in those observables remains. For $\sigma^{\text{DIS}}_{\gamma}(x)$, 28 data points are calculated, with $x=$ 0.0001, 0.0002, 0.0003, 0.0004, 0.0005, 0.0006, 0.0007, 0.0008, 0.0009, 0.001, 0.002, 0.003, 0.004, 0.005, 0.006, 0.007, 0.008, 0.009, 0.01, 0.02, 0.03, 0.04, 0.05, 0.06, 0.07, 0.08, 0.09, 0.1. For $\sigma^{\text{DIS}}_{Z}(x)$, 16 points are calculated, with $x=$ 0.01, 0.02, 0.03, 0.04, 0.05, 0.06, 0.07, 0.08, 0.09, 0.1, 0.15, 0.2, 0.25, 0.3, 0.35, 0.4.
Similarly, we introduce the following cross section observables in the proton-proton collisions: $\sigma^{\text{pp}}_{Z}(x_{1},x_{2})$, $\sigma^{\text{pp}}_{W^{+}}(x_{1},x_{2})$ and $\sigma^{\text{pp}}_{W^{-}}(x_{1},x_{2})$ as:
$\sigma^{\text{pp}}_{Z}(x_{1},x_{2})$ $=$ $\left[u_{V}(x_{1})+\bar{u}(x_{1})\right]\bar{u}(x_{2})+\left[u_{V}(x_{2})+\bar{u}(x_{2})\right]\bar{u}(x_{1})$
$+1.2\left[d_{V}(x_{1})+\bar{d}(x_{1})\right]\bar{d}(x_{2})$
$+1.2\left[d_{V}(x_{2})+\bar{d}(x_{2})\right]\bar{d}(x_{1})$
$+1.2\times 2\times s(x_{1})s(x_{2})$
$\sigma^{\text{pp}}_{W^{+}}(x_{1},x_{2})$ $=$ $0.9\left[(u_{V}(x_{1})+\bar{u}(x_{1}))\bar{d}(x_{2})\right]$
$+0.9\left[(u_{V}(x_{2})+\bar{u}(x_{2}))\bar{d}(x_{1})\right]$
$+0.1\left[(u_{V}(x_{1})+\bar{u}(x_{1}))s(x_{2})\right]$
$+0.1\left[(u_{V}(x_{2})+\bar{u}(x_{2}))s(x_{1})\right]$
$\sigma^{\text{pp}}_{W^{-}}(x_{1},x_{2})$ $=$ $0.9\left[(d_{V}(x_{1})+\bar{d}(x_{1}))\bar{u}(x_{2})\right]$ (19)
$+0.9\left[(d_{V}(x_{2})+\bar{d}(x_{2}))\bar{u}(x_{1})\right]$
$+0.1\bar{u}(x_{1})s(x_{2})+0.1\bar{u}(x_{2})s(x_{1})$
They approximately represent the vector boson productions of $pp(q\bar{q})\rightarrow Z/\gamma^{*}\rightarrow\ell^{+}\ell^{-}$ and $pp(q_{i}\bar{q}_{j})\rightarrow W^{\pm}\rightarrow\ell^{\pm}+\nu$. The momentum fractions $x_{1}$ and $x_{2}$ of the two initial quarks are connected by
$x_{1,2}=\frac{\sqrt{M^{2}+Q^{2}_{T}}}{\sqrt{s}}e^{\pm Y}$
where $M$, $Q_{T}$ and $Y$ are the invariant mass, transverse momentum and rapidity of the bosons, respectively. In this work, $M$ is set to 90 GeV, which is near the pole of the $W$ and $Z$ boson masses. Compared to $M$, the transverse momentum of the single vector boson production is usually small. Therefore $Q_{T}$ is ignored in this work. $\sqrt{s}$ is set to 13 TeV. To gain information for different $x$ values, the cross section observables are calculated with different $|Y|$ values of 0.1, 0.3, 0.5, 0.7, 0.9, 1.1, 1.3, 1.5, 1.7, 1.9, 2.1, 2.4, 2.8, 3.2, 4.0. The corresponding $x$ values range from 0.0001 to 0.38. In summary, 149 pseudo data points are introduced, corresponding to the DIS and Drell-Yan processes that are actually used in the PDF global analysis. For each data point, we assign a relative uncertainty of $3\%$. This is a precision comparable to the HERA measurements and the LHC Drell-Yan measurements. For experimental results other than HERA and LHC, such as earlier DIS and semi-inclusive DIS measurements, their precisions are much lower. The experiments actually used in the PDF global analysis are summarized in Ref. [2]. Although these pseudo observables are only rough estimations rather than real calculations, they reflect the complexcity of using multiple data inputs to determine the PDF shape parameters.
With the constructed PDFs, the pseudo observables and the assigned uncertainty, the Hessian matrix can be calculated accordingly. Since there are 24 free parameters in this test, the same number of eigenvectors and corresponding eigenvalues are acquired by diagonalizing the Hessian matrix. The eigenvalues in this test are shown in Fig. 2. As explained in Ref. [1], the eigenvalues in logarithmic scale rank as a line, reflecting the sensitivities of the input data. An eigenvector with large eigenvalue usually indicates that the information related to this eigenvector is well constrained by the input data. Otherwise, the input data has poor sensitivity to this eigenvector, leading to a small eigenvalue. Although data points in this analysis are much more than the free parameters being determined and the assigned uncertainty of $3\%$ is quite small, there is still some PDF information for certain $x$ regions and parton flavors, for which no pseudo observable gives good sensitivity. Therefore, some eigenvalues are small compared to others. This also happens in a real PDF global analysis [2, 3, 7].
Figure 2: 24 eigenvalues calculated from the Hessian matrix in the pseudo global analysis.
A simple test can numerically show whether the data inputs in a global fit provide a strong constraint to a particular eigenvector, as to be illustrated below. Once the Hessian matrix is diagonalized and the orthogonal displacements $\{z_{i}\}$ are acquired, one can look into the $\chi^{2}$-$z_{i}$ relationship. Usually, if the eigenvector represented by $z_{i}$ is well constrained, $\chi^{2}$ is expected to be an ideal quadratic function of $z_{i}$, giving $\chi^{2}(z_{i}=\pm n;z_{j}=0$ for $j\neq i)=n^{2}$. In Fig. 3, the $\chi^{2}$-$z_{i}$ relationship for $i=22$ and $i=6$ are shown and compared, as an example. For $i=22$, the eigenvector has an eigenvalue at $\mathcal{O}(10^{4})$, thus its $\chi^{2}$-$z_{i}$ relationship follows a good quadratic approximation. For $i=6$ however, the eigenvalue is at $\mathcal{O}(1)$, resulting in a non-symmetric $\chi^{2}$-$z_{i}$ relationship. When the eigenvalue is small, the normalization factor $s_{k}\approx\sqrt{2/\epsilon_{k}}$ in Eq. (II) is enlarged. Therefore, any bias, such as inaccuracy in numerical calculations, would be significantly amplified. As a result, $z_{i}=\pm 1$ does not always yield $\Delta\chi^{2}=\chi^{2}-\chi^{2}_{\text{min}}=1$, especially for those poorly constrained eigenvectors. In conclusion, when the $\chi^{2}$-$z_{i}$ relationship does not follow a good quadratic approximation, the 1 standard deviation cannot be represented by $z_{i}=\pm 1$. Therefore, in this work, the one standard deviation in the uncertainty estimation of an observable is calculated with the $z_{i}$ value which yields $\Delta\chi^{2}=\chi^{2}-\chi^{2}_{\text{min}}=1$.
Figure 3: $\chi^{2}$-$z_{i}$ relationship for $i=22$ (demonstrated in the right vertical axis) and $i=6$ (demonstrated in the left vertical axis) eigenvectors.
Following the method described in the previous section, the first-order and second-order PDF error sets are acquired according to Eq. (9) and Eq. (11), respectively. The uncertainty of any physical observable $\mathcal{O}$ can be estimated accordingly. To numerically study the difference between the original Hessian method and the improved Hessian method, we compare the relative uncertainties on $xf_{q}(x)$ for $u_{V}$, $d_{V}$, $\bar{u}$, $\bar{d}$ and $s$, which are shown in Fig. 4. The improved Hessian method uncertainties, in which the second-order terms are included, are in general larger than the first-order calculations. Furthermore, the difference between the two uncertainties generally increases in the small and large $x$ region where the PDF uncertainty is larger. However, due to the complexity of the mixture of the parton information in multiple experimental observables, the uncertainties contributed from the second-order terms can still be relatively sizable, as compared to the first-order terms, even when the PDF uncertainty is small, i.e. when the PDFs are better constrained.
Figure 4: Relative uncertainties on $xf_{u_{V}}(x)$, $xf_{d_{V}}(x)$, $xf_{\bar{u}}(x)$, $xf_{\bar{d}}(x)$ and $xf_{s}(x)$, calculated by including only the first-order derivatives (dashed magenta curves) and with the addition of the second-order derivatives (solid red curves), respectively. The lower panel in each plot gives the ratio of the two uncertainties.
IV Pseudo global analysis with new data
The eigenvectors in a global analysis are determined from a particular set of data inputs. Therefore, both the first-order and second-order uncertainty estimations depend on the eigenvectors. To further test the improved Hessian method when the eigenvectors change, we introduce new experimental observables into the global fit, which are the new proton structure parameters $P_{u}$ and $P_{d}$ factorized from the forward-backward asymmetry of the Drell-Yan process [8], and the boost asymmetry $A_{\text{boost}}$ in the diboson process [9]. These asymmetry variables were recently proposed at hadron colliders which can provide unique information on the proton structure. The $P_{u}$ and $P_{d}$ parameters factorize the proton structure information in $pp(u\bar{u})\rightarrow Z/\gamma^{*}\rightarrow\ell^{+}\ell^{-}$ and $pp(d\bar{d})\rightarrow Z/\gamma^{*}\rightarrow\ell^{+}\ell^{-}$ events. The two light quark initial states can be decoupled because the electroweak couplings of $u$-$\bar{u}$-$Z$ and $d$-$\bar{d}$-$Z$ are different. As a result, $P_{u}$ and $P_{d}$ separately provide the $u$ and $d$ quark information, which are mixed and indistinguishable in the measurement of the total cross section $\sigma^{pp}_{Z}$. $A_{\text{boost}}$ is the asymmetry defined in the diboson production [9]. For example, in the $pp(u\bar{d})\rightarrow W^{+}\gamma\rightarrow\ell^{+}\nu+\gamma$ process, the $W^{+}$ boson predominantly couples to the $\bar{d}$ quark while the photon couples to the $u$ quark. Consequently, the boost asymmetry $A_{\rm boost}^{W^{+}\gamma}$ has a positive large value, namely $\gamma$ is more boosted than the $W^{+}$ boson, because the initial state parton lunminosity $u(x_{1}){\bar{d}}(x_{2})$ is larger than $u(x_{2}){\bar{d}}(x_{1})$ for $x_{1}>x_{2}$, etc. Therefore, the relative asymmetry between the initial states $q_{i}(u_{1})\bar{d}(x_{2})$ and $u(x_{2})\bar{d}(x_{1})$ ($x_{1}>x_{2}$), which arises from the nature of the valence and sea quark differences, can be measured by comparing the energy or rapidity of the two bosons.
According to Ref. [8] and Ref. [9], the new pseudo observables can be constructed as:
$P_{u}(x_{1},x_{2})$ $=$ $\left[u_{V}(x_{1})+\bar{u}(x_{1})\right]\bar{u}(x_{2})-\left[u_{V}(x_{2})+\bar{u}(x_{2})\right]\bar{u}(x_{1})$
$P_{d}(x_{1},x_{2})$ $=$ $\left[d_{V}(x_{1})+\bar{d}(x_{1})\right]\bar{d}(x_{2})-\left[d_{V}(x_{2})+\bar{d}(x_{2})\right]\bar{d}(x_{1})$
$A^{W^{+}\gamma}_{\text{boost}}(x_{1},x_{2})$ $=$ $\left[u_{V}(x_{1})+\bar{u}(x_{1})\right]\bar{d}(x_{2})-\left[u_{V}(x_{2})+\bar{u}(x_{2})\right]\bar{d}(x_{1})$
$A^{W^{-}\gamma}_{\text{boost}}(x_{1},x_{2})$ $=$ $\left[d_{V}(x_{1})+\bar{d}(x_{1})\right]\bar{u}(x_{2})-\left[d_{V}(x_{2})+\bar{d}(x_{2})\right]\bar{u}(x_{1})$
To gain information on the $x$-dependence of the PDFs, the above pseudo observables are calculated with the $|Y|$ values the same as the previous settings. The relative uncertainty of each data point is also assigned as $3\%$. We repeat the calculations of the Hessian matrix now, using both the cross section observables and the new asymmetry observables; so we acquire the eigenvectors, eigenvalues, and the new orthogonal parameters $\{z_{i}\}$ accordingly. The new eigenvalues are shown in Fig. 5. With more data constraints introduced, the eigenvalues are in general larger. In Fig. 6, $\Delta\chi^{2}=\chi^{2}(z_{i}=1)-\chi^{2}_{\text{min}}$ is shown both in the previous pseudo global fit and the new fit. As discussed, $\Delta\chi^{2}=1$ indicates a good quadratic approximation, so that the corresponding eigenvector is well constrained by data. Therefore, more eigenvectors have $\Delta\chi^{2}=1$ or close to 1 in the new global fit, as expected.
Figure 5: 24 eigenvalues in the pseudo global analysis with new asymmetry observables included.
Figure 6: $\Delta\chi^{2}=\chi^{2}(z_{i}=1)-\chi^{2}_{\text{min}}$ in the pseudo global fit with new asymmetry observables included, compared with the previous global fit with only cross section observables.
The relative uncertainties on $xf_{q}(x)$ for $u_{V}$, $d_{V}$, $\bar{u}$, $\bar{d}$ and $s$ are shown in Fig. 7. As the constraints from the input data are strengthened, the estimated uncertainties are reduced with respect to the previous global fit in Fig. 4. In principle, when the uncertainties of individual eigenvectors decrease, the second-order uncertainties in Eq. (1) become less important with respect to the first-order ones. In Appendix B, we give another pseudo global fit in which the relative uncertainty of each data point is reduced to $0.1\%$. With such a high precision, there is indeed no difference between the first-order and second-order calculation, as expected. In the present section, although the new asymmetry observables provide stronger constraints, their precisions are not high enough to reduce the second-order terms to a negligible level. The ratios between the first-order and second-order uncertainties in Fig. 7 can become even larger than the previous global fit in Fig. 4, due to the complexity of the multiple data fit. This reflects the fact that the uncertainty estimation in a PDF global fit depends highly on the data inputs. Therefore, the second-order calculation is not only a necessary procedure to get a better uncertainty estimation, but also a sufficient way to evaluate the power of the data constraints in a global fit. When the difference between the first-order and second-order uncertainties is small, it indicates a strong data constraint so that the output PDF is less dependent on various assumptions and the choice of non-perturbative functions.
Figure 7: Relative uncertainty on $xf_{u_{V}}(x)$, $xf_{d_{V}}(x)$, $xf_{\bar{u}}(x)$, $xf_{\bar{d}}(x)$ and $xf_{s}(x)$. The new asymmetry observables $P_{u}$, $P_{d}$, and $A^{W^{\pm}\gamma}_{\text{boost}}$ are included in the global analysis.
V Conclusion
In this work, we propose an improved Hessian method in the PDF global analysis, which estimates the uncertainty with the inclusion of the complete set of second-order derivatives. A numerical test has been designed, in which a set of PDFs with multiple parameters are fitted to pseudo experimental observables. The test shows that when the precision of the input data used in a global analysis is limited, the improved Hessian method would give different uncertainties compared to the original method. Since the difference in the uncertainty can be reduced as the data constraints become stronger, one can tell whether a PDF global analysis is experimentally well constrained and less model-dependent, by comparing the uncertainties of the two methods.
VI Acknowledgement
Acknowledgements.
This work was supported by the National Natural Science Foundation of China under Grant No. 11721505, 11875245, 12061141005 and 12105275, and supported by the “USTC Research Funds of the Double First-Class Initiative”. This work was also supported by the U. S. National Science Foundation under Grant No. PHY-2310291.
Appendix A Appendix A: PDF-induced uncertainty on physical observables
In this section, we give detailed calculation of Eq. (14). Rewrite Eq. (1) as:
$\mathcal{O}(z_{i})$ $=$ $\mathcal{O}(z^{0}_{i})+\sum_{i}A_{i}+\frac{1}{2}\sum_{i\neq j}B_{ij}$
$A_{i}$ $=$ $\frac{\partial\mathcal{O}}{\partial z_{i}}z_{i}+\frac{1}{2}\frac{\partial^{2}\mathcal{O}}{\partial z^{2}_{i}}z^{2}_{i}$
$B_{ij}$ $=$ $\frac{\partial^{2}\mathcal{O}}{\partial z_{i}\partial z_{j}}z_{i}z_{j}$ (21)
The standard deviation of $A_{i}$ can be calculated as:
$\text{V}[A_{i}]$ $=$ $\text{E}[A^{2}_{i}]-\text{E}[A_{i}]^{2}$ (22)
$=$ $\text{E}\left[\left(\frac{\partial\mathcal{O}}{\partial z_{i}}\right)^{2}z^{2}_{i}\right]+\text{E}\left[\frac{\partial\mathcal{O}}{\partial z_{i}}\frac{\partial^{2}\mathcal{O}}{\partial z^{2}_{i}}z^{3}_{i}\right]$
$+\text{E}\left[\left(\frac{1}{2}\frac{\partial^{2}\mathcal{O}}{\partial z^{2}_{i}}\right)^{2}z^{4}_{i}\right]-\text{E}\left[A_{i}\right]^{2}$
where $\text{V}[X]$ and $\text{E}[X]$ corresponds to the standard deviation and mathematical expectation. Note that $\text{E}[z^{2}_{i}]=1$, $\text{E}[z^{3}_{i}]=0$ and $\text{E}[z^{4}_{i}]=3$ for the gaussian distribution $z_{i}$. Therefore, we have
$\text{V}[A_{i}]=\left(\frac{\partial\mathcal{O}}{\partial z_{i}}\right)^{2}+\frac{1}{2}\left(\frac{\partial^{2}\mathcal{O}}{\partial z^{2}_{i}}\right)^{2}.$ (23)
Similarly, the standard deviation of $B_{ij}$ is:
$\text{V}[B_{ij}]$ $=$ $\text{E}\left[B^{2}_{ij}\right]-\text{E}[B_{ij}]^{2}=\left(\frac{\partial^{2}\mathcal{O}}{\partial z_{i}z_{j}}\right)^{2}$ (24)
Given the fact that the $A_{i}$ and $B_{ij}$ terms are uncorrelated since they contain independent gaussian distributions, the standard deviation of $\mathcal{O}$ is simply the sum of $\text{V}[A_{i}]$ and $\text{V}[B_{ij}]$, which are positive values. Consequently, the inclusion of second-order derivative terms always increases the size of the PDF-induced uncertainty in the physical observable. Furthermore, including only the second-order diagonal terms generally underestimates the uncertainty.
Appendix B Appendix B: A pseudo global fit with high precision data
A pseudo global fit with high precision data is performed in this section. 9 data variables are used, including $\sigma^{\text{DIS}}_{\gamma}$, $\sigma^{\text{DIS}}_{Z}$, $\sigma^{pp}_{Z}$, $\sigma^{pp}_{W^{+}}$, $\sigma^{pp}_{W^{-}}$, $P_{u}$, $P_{d}$, $A^{W^{+}\gamma}_{\text{boost}}$ and $A^{W^{-}\gamma}_{\text{boost}}$. For each data point corresponding to a particular $x$ value, the relative uncertainty is set to $0.1\%$. With sych high precision data used in the global fit, the uncertainty of the second-order derivative calculation is consistent with the first-order calculation. The result is shown in Fig. 8.
Figure 8: Relative uncertainty on $xf_{u_{V}}(x)$, $xf_{d_{V}}(x)$, $xf_{\bar{u}}(x)$, $xf_{\bar{d}}(x)$ and $xf_{s}(x)$. The new asymmetry observables $P_{u}$, $P_{d}$, and $A^{W^{\pm}\gamma}_{\text{boost}}$ are included in the global analysis. Relative uncertainty of each data point is improved to $0.1\%$.

### v2

@@
 Abstract
-The Hessian method is widely used in the global analysis of parton distribution functions (PDFs).
-It uses a set of orthogonal eigenvectors to calculate the first-order derivative terms of a physics observable, so that the leading order PDF-induced uncertainty can be estimated accordingly.
-In this article, we report an improved Hessian method which introduces the second-order derivatives in the uncertainty estimation.
-To estimate the effect of the second-order uncertainties, a pseudo global analysis is designed.
-We also find that the PDF-induced uncertainties on physics variables can be significantly different by taking the second-order effects into account, particularly when the precisions of experimental inputs used in the PDF global analysis are limited.
-I Introduction
-In the past decades, a global analysis method based on quantum chromo-dynamics (QCD) has been established to study the proton structure.
-It sufficiently analyzes comprehensive data results from the deep inelastic scattering experiments with the energy scale of a few GeV, to the processes at hadron colliders at $\mathcal{O}(100)$ GeV.
-In practice, the global analysis is required to not only give the central predictions on parton densities corresponding to the best fit of the data, but also provide a way that the uncertainties can be easily calculated and extrapolated for any particular physics observable.
-The Hessian method [1] is one of the most widely used strategies to estimate the PDF uncertainties, such as in the latest CT18NNLO and MSHT20 PDFs [2, 3] which use both the DIS and hadron collider data.
+The Hessian method is widely applied in the global analysis of parton distribution functions (PDFs), which uses a set of orthogonal eigenvectors to give predictions of a physical observable.
+Its uncertainty is estimated based on the assumption that all physical observables can be approximately expressed as linear functions of the non-perturbative parameters in PDF.
+In this article, we report an improved Hessian method which takes the non-linear effects into account in the uncertainty estimation.
+A pseudo global analysis is designed to numerically test the new method.
+The non-linear uncertainties can significantly enlarge the original linear ones.
+Such uncertainties can be reduced with high precision data introduced into the global analysis.
+However, the non-linear effect could still be sizable corresponding to the current precision of the experimental measurements.
+I I. Introduction
+In the past decades, a global analysis method based on quantum chromodynamics (QCD) has been established to study the proton structure at high energy scale.
+It sufficiently analyzes comprehensive data results from the deep inelastic scattering (DIS) experiments with the energy scale of a few GeV, to the processes at hadron colliders at $\mathcal{O}(100)$ GeV.
+The global analysis is expected to not only give the central predictions on parton densities corresponding to the best fit of the data, but also provide a way that the uncertainties can be easily calculated and extrapolated for any particular physical observable.
+The Hessian method [1] is one of the most widely used strategies to estimate the PDF uncertainties, such as in the latest CT18NNLO and MSHT20 PDFs [2, 3].
 In this approach, a group of error PDF sets are provided, which represent the possible variations on the parameters of the non-perturbative functions of the quarks and gluons.
-These error sets correspond to orthogonal eigenvectors in the original parameter space, so that the uncertainty of a particular physics observable can be calculated using the error sets without complicated analysis on the correlations between the original parameters.
-However, in the current Hessian method, only the leading order terms in the expansion with respect to a physics observable of those eigenvectors are taken into account to estimate the PDF-induced errors.
-As more data inputs are included in the global analysis, and more parameters are used in the non-perturbative parameterization of the PDFs, the higher order terms in the expansion of a physics observable may become more important.
-This has been mentioned in a few previous studies.
-For example, in the Error PDF Updating Method Package [4], contributions from the second-order terms in the expansion of a single eigenvector, appearing as the diagonal terms in the Hessian matrix as to be discussed later, were estimated based on the CT14 PDFs [5].
+These error sets correspond to orthogonal eigenvectors in the original parameter space, so that the uncertainty of a particular physical observable can be calculated using the error sets without complicated analysis on the correlations between the original parameters.
+In the current Hessian method, the uncertainty estimation is established as a linear approximation.
+Firstly, the orthogonal eigenvectors and their corresponding uncertainties are determined by assuming the physical observables of those data used in the global analysis can be expressed as linear functions of the non-perturbative PDF parameters.
+Secondly, when computing the PDF-induced uncertainty on a given physical observable using the eigenvectors, the observable is also assumed as linear function of the eigenvectors.
+The non-linear effects, appearing as the higher order derivatives in the expansion of physical observables as functions of PDF parameters or eigenvectors, are ignored.
+Such effect could be significant in high energy interactions, such as the LHC which provides precise experimental results in today’s PDF global analysis.
+A few previous studies have been performed on the non-linearities in the global analysis.
+For example, Ref.
+[4] discussed contributions from the second-order terms based on the CT14 PDFs [5], but only for those appearing as the diagonal elements in the Hessian matrix (as to be introduced later).
 Based on the same PDFs, it was found later in Ref.
-[6] that the higher order terms of a single eigenvector can be sizable in the analysis of the CMS di-jet measurement.
-However, the off-diagonal terms corresponding to the expansion of multiple eigenvectors cannot be faithfully tested in the above studies since CT14 did not provide the needed error sets for this kind of tasks.
-In this work, we introduce an improved Hessian method, which generated the PDF error sets allowing a complete second-order calculation of the uncertainty of a physics observable.
+[6] that the diagonal higher order terms could be sizable in the analysis of the CMS di-jet measurement.
+In this work, we introduce an improved Hessian method, in which the non-linear effects are estimated as additional uncertainties.
 To make a detailed comparison between the improved method and the original one, a pseudo global fit is studied.
-As will be demonstrated in this test, when including the second-order terms, the uncertainty on parton densities could be noticeably different from the original leading order estimation.
+As will be demonstrated in this test, the uncertainty on parton densities could be noticeably different from the original linear estimation.
 Such difference can be reduced as the precision of the input data in the global analysis improves.
-However, the required precision is beyond many experimental inputs included in the present PDF global analysis, especially from those old DIS and fixed-target Drell-Yan data [2].
-Hence, it would be important to consider the higher order terms in the uncertainty estimation of the PDF global analysis.
-II Method
-The Hessian method [1] is designed based on the following expansion:
-$\mathcal{O}(z_{i})=\mathcal{O}(z^{0}_{i})+\sum_{i}\frac{\partial\mathcal{O}}{\partial z_{i}}\Big{|}_{z^{0}_{i}}\text{d}z_{i}+\frac{1}{2}\sum_{i,j}\frac{\partial^{2}\mathcal{O}}{\partial z_{i}\partial z_{j}}\Big{|}_{z^{0}_{i},z^{0}_{j}}\text{d}z_{i}\text{d}z_{j}+\cdot\cdot\cdot$ (1)
-where $\mathcal{O}$ is a physical observable whose theory prediction depends on PDFs.
-$\{z_{i}\}$ represent the parameters in the PDFs, and $\{z^{0}_{i}\}$ are the corresponding central values determined as the best fit to the data inputs.
-When $z_{i}$ is independent of $z_{j}$, for $i\neq j$, the leading order uncertainty on $\mathcal{O}$ can simply be calculated as:
-$\delta\mathcal{O}\approx\sqrt{\sum_{i}\left(\frac{\mathcal{O}(z^{0}_{i}+\delta z_{i})-\mathcal{O}(z^{0}_{i}-\delta z_{i})}{2}\right)^{2}}$ (2)
-where $\delta z_{i}$ is the uncertainty on $z_{i}$.
-$\mathcal{O}(z^{0}_{i}\pm\delta z_{i})$ are the calculations of $\mathcal{O}$ where the value of the $i$-th parameter is shifted by ±$\pm\delta z_{i}$, while all other parameters $z_{j}(j\neq i)$ are fixed at their central values $z^{0}_{j}$.
-To enable the calculation, the global analysis provides a central PDF set $S_{0}=S(z^{0}_{i})$ which can be used to calculate the central prediction of $\mathcal{O}(S_{0})$, and a group of error PDF sets $S_{i\pm}=S(z^{0}_{i}\pm\delta z_{i};z_{j}=z^{0}_{j})$ for $j\neq i$.
-In this strategy, the first-order derivatives are represented by
-$\frac{\partial\mathcal{O}}{\partial z_{i}}\Big{|}_{z^{0}_{i}}=\frac{\mathcal{O}(z^{0}_{i}+\delta z_{i})-\mathcal{O}(z^{0}_{i}-\delta z_{i})}{2\delta z_{i}},$ (3)
-while the higher order derivatives $\partial^{2}\mathcal{O}/\partial z_{i}\partial z_{j}$ are ignored.
-As discussed above, Ref.
-[4] and Ref.
-[6] studied effects on $\partial^{2}\mathcal{O}/\partial z^{2}_{i}$.
-We first review how to obtain the orthogonal eigenvectors $\{v_{i}\}$, with $\{z_{i}\}$ being the displacement (in the space of $\{z_{i}\}$) along the direction of $\{v_{i}\}$, by the Hessian method.
-The non-perturbative functions which describe the parton densities at the scale of $Q_{0}$, taken from the CT18NNLO PDF as an example, are written in the formalism as:
+However, the needed precision to reduce the non-linear uncertainties to negligible level is far beyond the current precision of experimental observations.
+The article will be arranged as following: In section II, we review the general strategy of Hessian method; In section III, we discuss the non-linear effects in the global analysis; In section IV, we introduce the improved method to estimate the non-linear uncertainties; In section V, we perform a pseudo global analysis study to numerically test the improved method, and summarize the conclusion of this article.
+II II. The original Hessian method
+In this section, we review the original Hessian method [1] and its linear uncertainty estimation.
+In Hessian method, a series of non-perturbative functions are introduced to describe the parton densities at the scale of $Q_{0}$.
+Taking from the CT18NNLO PDF as an example, the functions are written in the formalism as:
 $f_{q}(x,Q_{0})=a_{0}x^{a_{1}-1}(1-x)^{a_{2}}P_{q}(x;a_{3},a_{4},\cdot\cdot\cdot)$ (1)
@@
 $\{a_{i}\}$ are the original parameters being determined in the global fit, by minimizing the $\chi^{2}$ defined as:
-$\chi^{2}=\sum_{\alpha}\frac{[D_{\alpha}-T_{\alpha}(a_{i})]^{2}}{\sigma^{2}_{\alpha}}$ (5)
+$\chi^{2}=\sum_{\alpha}\frac{[D_{\alpha}-T_{\alpha}(\{a_{i}\})]^{2}}{\sigma^{2}_{\alpha}}$ (2)
 where $D_{\alpha}$ and $\sigma_{\alpha}$ are the experimental inputs and their corresponding uncertainties.
 $T_{\alpha}(a_{i})$ is the theory prediction of $D_{\alpha}$, depending on the PDF shape parameters $\{a_{i}\}$.
-In practice, $\chi^{2}_{\text{total}}$ is more complicated, which may include correlated systematic uncertainties.
-Furthermore, to deal with the potential biases of the theory calculation and the experimental measurements, possible inconsistency between data, and the uncertainty due to the choice of the non-perturbative parameterizations, some dynamical and/or global tolerances are also introduced to define the errors of the fitted PDF parameters, such as in the CT18 and MSHT20 [2, 3].
-For simplicity, we ignore all these variants in this work as they will not strongly alter our conclusions.
-In this work, we treat the pseudo-PDF uncertainty to be uncorrelated and purely statistical.
 Given the fact that the best fit values of $\{a_{i}\}$ appear near the minimal $\chi^{2}$, we have:
@@
 The Hessian matrix is then defined as $H_{ij}=\partial^{2}\chi^{2}/\partial a_{i}\partial a_{j}$.
-Note that the original parameters $a_{i}$ can be correlated in the global fit because experimental observables usually relate to multiple PDF shape parameters.
+Note that $\{a_{i}\}$ are quite independent as free parameters in the non-perturbative functions, but their uncertainties in a global analysis can be correlated because experimental observables usually relate to multiple PDF shape parameters.
+The correlation depends on the particular set of experimental inputs used in a global fit.
 One can convert $\{a_{i}\}$ to another set of orthogonal parameters $\{z_{i}\}$ by diagonalizing $H_{ij}$:
@@
 $a_{i}-a^{0}_{i}$ $=$ $\sum_{k}v_{ik}s_{k}z_{k},$ (4)
-where the scaling factor $s_{i}$ is introduced to normalize $z_{i}$. Usually, $s_{k}$ is chosen to make
+where the scaling factors $\{s_{i}\}$ are introduced to normalize $\{z_{i}\}$. Usually, $\{s_{i}\}$ are chosen to make
 $\Delta\chi^{2}=\chi^{2}-\chi^{2}_{\text{min}}=\sum_{i}z^{2}_{i},$ (5)
 so that the best fit of $\chi^{2}$ (denoted as $\chi^{2}_{\text{min}}$) corresponds to $\{z_{i}=0\}$, and the one-standard deviation uncertainty in the direction of the $i$-th eigenvector corresponds to $\{z_{i}=\pm 1;z_{j}=0$ for $j\neq i\}$.
-By this requirement, $s_{i}$ are approximately equal to $\sqrt{2/\epsilon_{i}}$.
-The central PDF set and the error sets can be established.
-The central PDF set corresponds to the best fit parameters $S_{0}=S(a_{i}=a^{0}_{i})$, or $S_{0}=S(z_{i}=0)$.
-For the error PDF sets, we denote the ±$\pm 1$ standard variation in the direction of the $i$-th eigenvector as
-$S_{i\pm}=S(z_{i}=\pm 1;z_{j}=0(j\neq i)).$ (9)
-Hence, the derivative in Eq. (3) of any physical observable $\mathcal{O}$ can be calculated using the PDF error sets:
-$\frac{\partial\mathcal{O}}{\partial z_{i}}\Big{|}_{0}=\frac{\mathcal{O}(S_{i+})-\mathcal{O}(S_{i-})}{2},$ (10)
-so that the leading order uncertainty is acquired.
-$S_{0}$ and $S_{i\pm}$ are the PDF sets provided by the global analysis groups such as CTEQ-TEA and MSHT.
-They can be used to calculate the first-order derivatives in Eq.
-(1), but not those off-diagonal terms.
-Below, we show how to obtain the PDF error sets for calculating the full second-order derivatives in Eq. (1). Following the definition of the derivative, the needed error sets are:
+By this requirement, $\{s_{i}\}$ are approximately equal to $\sqrt{2/\epsilon_{i}}$.
+In practice, $\chi^{2}$ is more complicated, which may include correlated systematical uncertainties.
+Furthermore, the one-standard deviation uncertainty is usually enlarged by some dynamical and/or global tolerances to deal with the possible inconsistency between data, and to cover the uncertainty due to the choice of the non-perturbative parameterizations.
+For simplicity, we ignore all these variants in this work as they will not strongly alter our conclusions.
+For a given physical observable whose theory prediction depends on PDFs, it can be expanded as:
+$\mathcal{O}(\{z_{i}\})=\mathcal{O}(\{z^{0}_{i}\})+\sum_{i}\frac{\partial\mathcal{O}}{\partial z_{i}}\Big{|}_{z^{0}_{i}}\text{d}z_{i}+\frac{1}{2}\sum_{i,j}\frac{\partial^{2}\mathcal{O}}{\partial z_{i}\partial z_{j}}\Big{|}_{z^{0}_{i},z^{0}_{j}}\text{d}z_{i}\text{d}z_{j}+\cdot\cdot\cdot$ (6)
+As $\{z_{i}\}$ represent orthogonal eigenvectors, the leading order uncertainty on $\mathcal{O}$ can simply be calculated as:
+$\delta\mathcal{O}$ $\approx$ $\sqrt{\sum_{i}\left[\frac{\mathcal{O}(z^{0}_{i}+\delta z_{i})-\mathcal{O}(z^{0}_{i}-\delta z_{i})}{2\delta z_{i}}\right]^{2}}$ (7)
+where $\delta z_{i}=1$ is the uncertainty on $z_{i}$.
+$\mathcal{O}(z^{0}_{i}\pm\delta z_{i})$ is the calculation of $\mathcal{O}$ where the value of the $i$-th parameter is shifted by $\pm\delta z_{i}$, while all other parameters $z_{j}(j\neq i)$ are fixed at their central values $z^{0}_{j}$.
+To enable the calculation, the global analysis provides a central PDF set $S_{0}=S(\{z^{0}_{i}\})$ which can be used to calculate the central prediction of $\mathcal{O}(S_{0})$, and a group of error PDF sets $S_{i\pm}=S(z^{0}_{i}\pm\delta z_{i};z_{j}=z^{0}_{j})$ for $j\neq i$.
+Hence, the uncertainty in Eq.
+(7) can be calculated using the PDF error sets:
+$\delta\mathcal{O}=\sqrt{\sum_{i}\left(\frac{\mathcal{O}(S_{i+})-\mathcal{O}(S_{i-})}{2}\right)^{2}}$ (8)
+III III. Non-linear effects in global analysis
+The first linear assumption in the original Hessian method is introduced in the determination of the eigenvectors and their uncertainties.
+The orthogonal parameters $\{z_{i}\}$ are acquired using the Hessian matrix containing only the $\partial^{2}\chi^{2}/\partial a_{i}\partial a_{j}$ terms.
+The one standard deviation is defined as $\delta z_{i}=1$ corresponding to $\Delta\chi^{2}=1$.
+All these settings are based on the assumption that the physical observables of the data used in the global analysis can be written as the linear functions of the original PDF parameters $\{a_{i}\}$:
+$T_{\alpha}(\{a_{i}\})=\sum_{i}t^{i}_{\alpha}a_{i}$ (9)
+where $t^{i}_{\alpha}$ is the coefficient.
+In principle, Eq.
+(9) never holds true in global analysis.
+Firstly, the non-perturbative functions in Eq.
+(1) are naturally non-linear, e.g.
+the power terms $x^{1-a_{1}}$ and $(1-x)^{a_{2}}$ which are guided by the Regge theory [7] and the spectator counting rules [8] to describe the asymptotic behavior of $f_{q}(x)$ in the limits $x\rightarrow 0$ and 1.
+Secondly, the calculations on physical interactions are usually convolutions over $x$.
+Thirdly, for those collider and fixed target experiments, the calculations on the two-hadron initial state contain the product of the two parton densities.
+In practice, when data inputs used in a global analysis have high precisions, $\Delta\chi^{2}$ would increase rapidly as $\{a_{i}\}$ values running away from their best fitted values.
+Thus, all calculations can be done in small intervals around $\{a^{0}_{i}\}$ so that $T_{\alpha}(\{a_{i}\})$ can be numerically linear.
+In other words, the non-linearity depends on the precision of the experimental measurements in the PDF global analysis.
+In current PDFs, the experimental inputs provide good constraints in some $x$ regions and for the quark flavors that they are sensitive to.
+However, the data precisions in the foreseeable future would not be good enough to suppress the non-linear effect for all quark flavors and in a wide range of $x$.
+Another linear assumption is introduced in computing the PDF-induced uncertainties on physical observables using the eigenvectors.
+In the current Hessian method, the higher order contributions in Eq.
+(6) are ignored when extrapolating the uncertainties on eigenvectors to a given physical observable.
+The Hessian-type PDFs such as CT18 and MSHT20 provide only the error sets $S_{i\pm}$ corresponding to the complete differential of $z_{i}$, which allows calculation of $\partial\mathcal{O}/\partial z_{i}$, $\partial^{2}\mathcal{O}/\partial z^{2}_{i}$ or even higher order contributions as the diagonal elements.
+But these error sets cannot give estimations on the non-linear terms such as $\partial^{2}\mathcal{O}/\partial z_{i}\partial z_{j}$.
+The higher order contributions can be significant especially when the computed physical observable corresponds to $x$ region or quark flavor very differently from that of the data in the global fit.
+In CT18 [2], the uncertainties of the PDF error sets are quoted to cover the difference from varying the formalism of the non-perturbative functions of Eq.
+(1).
+By doing this, CT18 gives more conservative uncertainty estimation compared to MSHT20 [3] and NNPDF4 [10], which use similar data set in the global fit.
+The enlarged uncertainties reflect the non-linear effects in global analysis, because if the non-perturbative functions tested by CT18 were all in good linear approximation, they would lead to consistent predictions on any physical observable because they were connected by linear transformations in the PDF parameter space.
+In other words, it is, to some extent, similar to estimating the uncertainty due to the non-linearity with respect to the uncertainty due to the particular choice of non-perturbative functions used in the global fit.
+However, it is inevitable in the enumeration method that many functions being tested are known to have disadvantages in describing the data.
+Consequently, the uncertainty may be overestimated by quoting the imperfection of the excluded functions rather than the imperfection of the chosen function.
+On the other hand, the uncertainty may also be underestimated as there is always possible functions missed in the enumeration.
+Therefore, a new method to estimate the uncertainty of the chosen function itself would be important.
+IV IV. Improved Hessian method
+The improved method will be derived in this Section.
+Then, in Section V we will apply it to a pseudo analysis, to test how non-linear contributions can affect a global analysis.
+The pseudo-linear calculation will be compared to the nominal (un-improved) global fit.
+The strategy is as following:
+i) For a given non-perturbative function formalism such as in Eq.
+(1), perform the global analysis with the normal method, giving the best fitted values of $\{a_{i}=a^{0}_{i}\}$ and the central PDF set $S_{0}$.
+In this nominal fit, $T_{\alpha}(\{a_{i}\})$ are predicted with different $\{a_{i}\}$ inputs based on the conventional calculations.
+ii) For each physical observable, a set of linear factors are defined as:
+$K^{\alpha}_{i}$ $=$ $\frac{T^{i+}_{\alpha}-T^{i-}_{\alpha}}{2\epsilon_{i}}.$ (10)
+where $T^{i+}_{\alpha}=T_{\alpha}(\{a_{i}=a^{0}_{i}+\epsilon_{i};a_{j}=a^{0}_{j}\text{ for }i\neq j\})$ and $T^{i-}_{\alpha}=T_{\alpha}(\{a_{i}=a^{0}_{i}-\epsilon_{i};a_{j}=a^{0}_{j}\text{ for }i\neq j\})$.
+Apparently, $K^{\alpha}_{i}$ is an approximation of $\partial T_{\alpha}(\{a_{i}\})/\partial a_{i}$ around $\{a_{i}=a^{0}_{i}\}$ in the range of $(a_{i}-\epsilon_{i},a_{i}+\epsilon_{i})$.
+It represents the slope of the secant line, as shown in Fig.
+1.
+The $\epsilon_{i}$ parameter corresponds to the range of $a_{i}$ in which the linear derivation of Eq.
+(10) is assumed.
+Thus, $\epsilon_{i}$ can simply be set to the original uncertainty of the $a_{i}$ parameter given by the fit obtained form i).
+Figure 1: The secant line of $T_{\alpha}(a_{i};a_{j}=a^{0}_{j}\text{ for }j\neq i)$ around the best fitted value $a^{0}_{i}$.
+iii) A group of new observables can then be established as linear functions of $\{a_{i}\}$:
+$T^{(r)}_{\alpha}(\{a_{i}\})=T_{\alpha}(\{a^{r}_{i}\})+\sum_{i}K^{\alpha}_{i}\times(a_{i}-a^{r}_{i}),$ (11)
+where $\{a^{r}_{i}\}$, $r=0,1,\cdot\cdot\cdot,n$ represent a set of points in the parameter space of $\{a_{i}\}$, with $n$ as the number of the parameters.
+Values of $\{a^{r}_{i}\}$ also represent the range of the linear assumption, but in terms of $T_{\alpha}$.
+In order to achieve this, $\{a^{r}_{i}\}$ values can be set corresponding to $\{z_{r}=\Delta_{r};z_{i}=0\text{ for }i\neq r\}$.
+$\Delta_{r}$ is chosen to have $T_{\alpha}(\{a^{r}_{i}\})-T_{\alpha}(\{a^{0}_{i}\})$ consistent with $\sigma_{\alpha}$ for those data inputs sensitive to $z_{r}$.
+Note that when the eigenvector $z_{r}$ is constrained by multiple data inputs in the global fit, the uncertainty of $z_{r}$ ($\delta z_{r}$) will be smaller than $\Delta_{r}$, since $\delta z_{r}$ reflects the combination of the statistical power of those data inputs.
+Using Eq. (11), one can perform a linear fit, in which the best values of $\{a_{i}\}$ are determined by requiring the minimum of $\chi^{2}(T^{(r)})$ defined as:
+$\chi^{2}(T^{(r)})=\sum_{\alpha}\frac{[D_{\alpha}-T^{(r)}_{\alpha}(\{a_{i}\})]^{2}}{\sigma^{2}_{\alpha}}.$ (12)
+When Eq.
+(9) is in good approximation, the linear fit using Eq.
+(12) would give best fit values, denoted as $\{a^{r,0}_{i}\}$ for each $r$, consistent with the nominal fit $\{a^{0}_{i}\}$.
+Therefore, the difference between the two fits reflects the non-linearity of the physical observables $T_{\alpha}(\{a_{i}\})$, and can be quoted as additional uncertainties.
+To do this, an additional PDF error set is defined as $S_{E}(r)=S(\{a^{r,0}_{i}\})$ for each eigenvector.
+For a given physical observable $\mathcal{O}$, the non-linear uncertainty is defined as:
+$\delta\mathcal{O}_{E}=\sqrt{\sum_{r}\left[\mathcal{O}(S_{E}(r))-\mathcal{O}(S_{0})\right]^{2}}$ (13)
+It should be emphasized that Eq.
+(13) quantifies the impact on non-linearity when extrapolated to a physical observable.
+Such non-linearity introduces a bias in the analysis, which may in turn contribute to the PDF uncertainty.
+The notions of “bias” and “uncertainty” are not strictly equivalent, although they are likely to be of comparable size.
+A more detailed analysis would be required to establish their precise relationship in a specific case.
+$\delta\mathcal{O}_{E}$ is the uncertainty of assuming $T_{\alpha}(\{a_{i}\})$ as linear functions, which is independent of how to extrapolate the original experimental uncertainties $\sigma_{\alpha}$ to each eigenvector, and further to a physical observable $\mathcal{O}$.
+As discussed in the section II, the error sets $S_{i\pm}$ allow calculations only on the first-order derivatives in Eq.
+(6).
+Below, we show how to obtain the PDF error sets for calculating the full second-order derivatives.
+Following the definition of the derivative, the needed error sets are:
 $S^{i\pm}_{j\pm}=S(z_{i}=\pm 1,z_{j}=\pm 1;z_{k}=0\ \mathrm{for}\ k\neq i,j),$ (14)
-in which two displacements are varied simultaneously, so that the second-order derivatives in Eq. (3) can be calculated as:
+in which two displacements are varied simultaneously, so that the second-order derivatives can be calculated as:
 $\frac{\partial^{2}\mathcal{O}}{\partial z_{i}\partial z_{j}}$ $=$
 $\frac{1}{2}$ $\times$ $\left(\frac{\mathcal{O}(S^{i+}_{j+})-\mathcal{O}(S^{i-}_{j+})}{2}-\frac{\mathcal{O}(S^{i+}_{j-})-\mathcal{O}(S^{i-}_{j-})}{2}\right)$ (15)
-Note that for $i=j$, the second-order error sets are $S_{i++}=S(z_{i}=2;z_{k}=0$ for $k\neq i)$ and $S_{i--}=S(z_{i}=-2;z_{k}=0$ for $k\neq i)$. Consequently, we have
+For $i=j$, the second-order error sets can be calculated as $S_{i++}=S(z_{i}=2;z_{k}=0$ for $k\neq i)$ and $S_{i--}=S(z_{i}=-2;z_{k}=0$ for $k\neq i)$. Consequently, we have
 $\frac{\partial^{2}\mathcal{O}}{\partial z^{2}_{i}}$ $=$
@@
 Since the Hessian matrix comes from the derivatives of $\chi^{2}$, the second-order error sets can only be acquired together with the first-order error sets in a complete PDF global analysis.
-With full second-order derivatives included, the uncertainty on $\mathcal{O}$ is written as:
+With the full second-order derivatives included, the uncertainty on $\mathcal{O}$ is written as:
 $\left(\delta\mathcal{O}\right)^{2}=\sum_{i}\left(\frac{\partial\mathcal{O}}{\partial z_{i}}\right)^{2}+\frac{1}{2}\sum_{i}\left(\frac{\partial^{2}\mathcal{O}}{\partial z^{2}_{i}}\right)^{2}+\frac{1}{2}\sum_{i\neq j}\left(\frac{\partial^{2}\mathcal{O}}{\partial z_{i}\partial z_{j}}\right)^{2},$ (17)
 where the first, second and third terms correspond to the contributions from the first-order derivatives, the second-order diagonal derivatives, and the second-order off-diagonal derivatives in Eq.
-(1).
+(6), respectively.
 The process of proof of Eq.
@@
 (17).
-It is also convenient to calculate the correlation of the PDF-induced uncertainties between two physical observables $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$.
+It is also convenient for calculating the correlation of the PDF-induced uncertainties between two physical observables $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$.
 One can calculate $(\delta\mathcal{O}_{1})^{2}$, $(\delta\mathcal{O}_{2})^{2}$ and $(\delta(\mathcal{O}_{1}+\mathcal{O}_{2}))^{2}$ via either aforementioned method.
@@
 $\rho(\mathcal{O}_{1},\mathcal{O}_{2})=\frac{(\delta(\mathcal{O}_{1}+\mathcal{O}_{2}))^{2}-(\delta\mathcal{O}_{1})^{2}-(\delta\mathcal{O}_{2})^{2}}{2\delta\mathcal{O}_{1}\delta\mathcal{O}_{2}}$ (18)
-III Pseudo global analysis
+Note that both the bootstrap method and Eq.
+(17) assume $\{z_{i}\}$ have independent gaussian distributions.
+As will be discussed in the next section, it is equivalent to the linear assumptions of Eq.
+(9), of which the uncertainty has been estimated as $\delta\mathcal{O}_{E}$.
+In conclusion, the total uncertainty on a physical observable $\mathcal{O}$ is the quadratic combination of the original first-order uncertainties, the second-order uncertainties, and the non-linear uncertainty of $\delta\mathcal{O}_{E}$:
+$\left(\delta\mathcal{O}_{\text{total}}\right)^{2}=\frac{1}{4}\sum_{i}\left[\mathcal{O}(S_{i+})-\mathcal{O}(S_{i-})\right]^{2}$ (19)
+$+$ $\frac{1}{32}\sum_{i}\left[\mathcal{O}(S_{i++})+\mathcal{O}(S_{i--})-2\mathcal{O}(S_{0})\right]^{2}$
+$+$ $\frac{1}{32}\sum_{i\neq j}\left[\mathcal{O}(S^{i+}_{j+})-\mathcal{O}(S^{i-}_{j+})-\mathcal{O}(S^{i+}_{j-})+\mathcal{O}(S^{i-}_{j-})\right]^{2}$
+$+$ $\sum_{r}\left[\mathcal{O}(S_{E}(r))-\mathcal{O}(S_{0})\right]^{2}.$
+[…]
+V V. Pseudo global analysis
 In this section, we introduce a pseudo global analysis to numerically test the improved Hessian method.
@@
 $a_{3}$ 1.502 2.615 4 (fixed) 4 (fixed) 4(fixed)
-$a_{4}$0.147 1.828 0.618 0.292 0.466
+$a_{4}$ -0.147 1.828 0.618 0.292 0.466
 $a_{5}$ 1.671 2.721 0.195 0.647 0.466
-$a_{6}$ 0.871 0.474 0.225
-$a_{7}$ 0.267 0.741 0.225
-$a_{8}$ 0.733 1 (fixed) 1 (fixed)
+$a_{6}$ - - 0.871 0.474 0.225
+$a_{7}$ - - 0.267 0.741 0.225
+$a_{8}$ - - 0.733 1 (fixed) 1 (fixed)
 Table 1: Values of parameters in the non-perturbative functions, taken from CT18NNLO.
@@
 A group of pseudo observables are also constructed according to the PDFs in Eq.
-(III) and Eq.
-(III).
+(V) and Eq.
+(V).
 These observables, with uncertainties assigned, will be used as the data input in the pseudo global analysis.
@@
 The $q(x)$ functions are taken from Eq.
-(III) and Eq.
-(III), with $\{a_{i}\}$ values from Tab.
+(V) and Eq.
+(V), with $\{a_{i}\}$ values from Tab.
 1.
@@
 For $\sigma^{\text{DIS}}_{Z}(x)$, 16 points are calculated, with $x=$ 0.01, 0.02, 0.03, 0.04, 0.05, 0.06, 0.07, 0.08, 0.09, 0.1, 0.15, 0.2, 0.25, 0.3, 0.35, 0.4.
-Similarly, we introduce the following cross section observables in the proton-proton collisions: $\sigma^{\text{pp}}_{Z}(x_{1},x_{2})$, $\sigma^{\text{pp}}_{W^{+}}(x_{1},x_{2})$ and $\sigma^{\text{pp}}_{W^{-}}(x_{1},x_{2})$ as:
+Similarly, we introduce the following cross section observables in proton-proton collisions: $\sigma^{\text{pp}}_{Z}(x_{1},x_{2})$, $\sigma^{\text{pp}}_{W^{+}}(x_{1},x_{2})$ and $\sigma^{\text{pp}}_{W^{-}}(x_{1},x_{2})$ as:
 $\sigma^{\text{pp}}_{Z}(x_{1},x_{2})$ $=$ $\left[u_{V}(x_{1})+\bar{u}(x_{1})\right]\bar{u}(x_{2})+\left[u_{V}(x_{2})+\bar{u}(x_{2})\right]\bar{u}(x_{1})$
@@
 The corresponding $x$ values range from 0.0001 to 0.38.
+We also introduce a group of new experimental observables into the global fit, which are the proton structure parameters, $P_{u}$ and $P_{d}$, factorized from the forward-backward asymmetry of the Drell-Yan process [11], and the boost asymmetry $A_{\text{boost}}$ in the diboson process [12].
+These asymmetry variables were recently proposed at hadron colliders which can provide unique information on the proton structure.
+The $P_{u}$ and $P_{d}$ parameters factorize the proton structure information in $pp(u\bar{u})\rightarrow Z/\gamma^{*}\rightarrow\ell^{+}\ell^{-}$ and $pp(d\bar{d})\rightarrow Z/\gamma^{*}\rightarrow\ell^{+}\ell^{-}$ events.
+The two light quark initial states can be decoupled because the electroweak couplings of $u¯\bar{u}-$$Z$ and $d¯\bar{d}-$$Z$ are different.
+As a result, $P_{u}$ and $P_{d}$ separately provide the $u$ and $d$ quark information, which are mixed and indistinguishable in the measurement of the total cross section $\sigma^{pp}_{Z}$.
+$A_{\text{boost}}$ is the asymmetry defined in the diboson production [12].
+For example, in the $pp(u\bar{d})\rightarrow W^{+}\gamma\rightarrow\ell^{+}\nu+\gamma$ process, the $W^{+}$ boson predominantly couples to the $\bar{d}$ quark while the photon couples to the $u$ quark.
+Consequently, the boost asymmetry $A_{\rm boost}^{W^{+}\gamma}$ has a positive large value, namely $\gamma$ is more boosted than the $W^{+}$ boson, because the initial state parton lunminosity $u(x_{1}){\bar{d}}(x_{2})$ is larger than $u(x_{2}){\bar{d}}(x_{1})$ for $x_{1}>x_{2}$, etc.
+Therefore, the relative asymmetry between the initial-state parton luminosities $u(x_{1})\bar{d}(x_{2})$ and $u(x_{2})\bar{d}(x_{1})$ ($x_{1}>x_{2}$) can be measured by comparing the energy or rapidity of the two bosons.
+According to Ref. [11] and Ref. [12], the new pseudo observables can be constructed as:
+$P_{u}(x_{1},x_{2})$ $=$ $\left[u_{V}(x_{1})+\bar{u}(x_{1})\right]\bar{u}(x_{2})-\left[u_{V}(x_{2})+\bar{u}(x_{2})\right]\bar{u}(x_{1}),$
+$P_{d}(x_{1},x_{2})$ $=$ $\left[d_{V}(x_{1})+\bar{d}(x_{1})\right]\bar{d}(x_{2})-\left[d_{V}(x_{2})+\bar{d}(x_{2})\right]\bar{d}(x_{1}),$
+$A^{W^{+}\gamma}_{\text{boost}}(x_{1},x_{2})$ $=$ $\left[u_{V}(x_{1})+\bar{u}(x_{1})\right]\bar{d}(x_{2})-\left[u_{V}(x_{2})+\bar{u}(x_{2})\right]\bar{d}(x_{1}),$
+$A^{W^{-}\gamma}_{\text{boost}}(x_{1},x_{2})$ $=$ $\left[d_{V}(x_{1})+\bar{d}(x_{1})\right]\bar{u}(x_{2})-\left[d_{V}(x_{2})+\bar{d}(x_{2})\right]\bar{u}(x_{1}).$
 In summary, 149 pseudo data points are introduced, corresponding to the DIS and Drell-Yan processes that are actually used in the PDF global analysis.
 For each data point, we assign a relative uncertainty of $3\%$.
-This is a precision comparable to the HERA measurements and the LHC Drell-Yan measurements.
+This precision is comparable to the HERA measurements and the LHC Drell-Yan measurements.
 For experimental results other than HERA and LHC, such as earlier DIS and semi-inclusive DIS measurements, their precisions are much lower.
@@
 [2].
-Although these pseudo observables are only rough estimations rather than real calculations, they reflect the complexcity of using multiple data inputs to determine the PDF shape parameters.
+Although these pseudo observables are only rough estimations rather than real calculations, they reflect the complexity of using multiple data inputs to determine the PDF shape parameters.
 With the constructed PDFs, the pseudo observables and the assigned uncertainty, the Hessian matrix can be calculated accordingly.
@@
 The eigenvalues in this test are shown in Fig.
-2.
+3.
 As explained in Ref.
@@
 Otherwise, the input data has poor sensitivity to this eigenvector, leading to a small eigenvalue.
-Although data points in this analysis are much more than the free parameters being determined and the assigned uncertainty of $3\%$ is quite small, there is still some PDF information for certain $x$ regions and parton flavors, for which no pseudo observable gives good sensitivity.
+Although data points in this analysis are much more than the free parameters being determined and the assigned uncertainty of $3\%$ is quite small, there is still PDF information for certain $x$ regions and parton flavors, for which no pseudo observable gives direct sensitivity.
 Therefore, some eigenvalues are small compared to others.
@@
 Figure 3: 24 eigenvalues calculated from the Hessian matrix in the pseudo global analysis.
-A simple test can numerically show whether the data inputs in a global fit provide a strong constraint to a particular eigenvector, as to be illustrated below.
 Once the Hessian matrix is diagonalized and the orthogonal displacements $\{z_{i}\}$ are acquired, one can look into the $\chi^{2}$-$z_{i}$ relationship.
+It is a good numerical test of the linear assumptions in the global analysis.
 Usually, if the eigenvector represented by $z_{i}$ is well constrained, $\chi^{2}$ is expected to be an ideal quadratic function of $z_{i}$, giving $\chi^{2}(z_{i}=\pm n;z_{j}=0$ for $j\neq i)=n^{2}$.
+As discussed before, it also means the data inputs related to this eigenvector $z_{i}$ are precise enough so that the corresponding $T_{\alpha}$ are approximately linear functions of the original parameters $\{a_{i}\}$ within small intervals.
 In Fig.
-3, the $\chi^{2}$-$z_{i}$ relationship for $i=22$ and $i=6$ are shown and compared, as an example.
-For $i=22$, the eigenvector has an eigenvalue at $\mathcal{O}(10^{4})$, thus its $\chi^{2}$-$z_{i}$ relationship follows a good quadratic approximation.
-For $i=6$ however, the eigenvalue is at $\mathcal{O}(1)$, resulting in a non-symmetric $\chi^{2}$-$z_{i}$ relationship.
-When the eigenvalue is small, the normalization factor $s_{k}\approx\sqrt{2/\epsilon_{k}}$ in Eq.
-(II) is enlarged.
-Therefore, any bias, such as inaccuracy in numerical calculations, would be significantly amplified.
-As a result, $z_{i}=\pm 1$ does not always yield $\Delta\chi^{2}=\chi^{2}-\chi^{2}_{\text{min}}=1$, especially for those poorly constrained eigenvectors.
-In conclusion, when the $\chi^{2}$-$z_{i}$ relationship does not follow a good quadratic approximation, the 1 standard deviation cannot be represented by $z_{i}=\pm 1$.
-Therefore, in this work, the one standard deviation in the uncertainty estimation of an observable is calculated with the $z_{i}$ value which yields $\Delta\chi^{2}=\chi^{2}-\chi^{2}_{\text{min}}=1$.
+4, the $\chi^{2}$-$z_{i}$ relationship for $i=22$ and $i=6$ are shown and compared, as an example.
+$i=22$ is a typical eigenvector which has good linear approximation.
+Its eigenvalue is large at $\mathcal{O}(10^{4})$.
+The $\chi^{2}$-$z_{i}$ relationship follows a good quadratic approximation, so that $z_{i}=1$ yields $\Delta\chi^{2}=\chi^{2}-\chi^{2}_{\text{min}}=1$ corresponding to 1 standard deviation.
+For $i=6$ however, it is an eigenvector which is not well constrained by data, resulting in a non-symmetric $\chi^{2}$-$z_{i}$ relationship.
+In practice (also in this work), the 1 standard deviation in the uncertainty estimation is usually defined with the $\{z_{i}\}$ values which yields $\Delta\chi^{2}=\chi^{2}-\chi^{2}_{\text{min}}=1$.
 Figure 4: $\chi^{2}$-$z_{i}$ relationship for $i=22$ (demonstrated in the right vertical axis) and $i=6$ (demonstrated in the left vertical axis) eigenvectors.
-Following the method described in the previous section, the first-order and second-order PDF error sets are acquired according to Eq.
-(9) and Eq.
-(11), respectively.
-The uncertainty of any physical observable $\mathcal{O}$ can be estimated accordingly.
-To numerically study the difference between the original Hessian method and the improved Hessian method, we compare the relative uncertainties on $xf_{q}(x)$ for $u_{V}$, $d_{V}$, $\bar{u}$, $\bar{d}$ and $s$, which are shown in Fig.
-4.
-The improved Hessian method uncertainties, in which the second-order terms are included, are in general larger than the first-order calculations.
-Furthermore, the difference between the two uncertainties generally increases in the small and large $x$ region where the PDF uncertainty is larger.
-However, due to the complexity of the mixture of the parton information in multiple experimental observables, the uncertainties contributed from the second-order terms can still be relatively sizable, as compared to the first-order terms, even when the PDF uncertainty is small, i.e.
-when the PDFs are better constrained.
-Figure 4: Relative uncertainties on $xf_{u_{V}}(x)$, $xf_{d_{V}}(x)$, $xf_{\bar{u}}(x)$, $xf_{\bar{d}}(x)$ and $xf_{s}(x)$, calculated by including only the first-order derivatives (dashed magenta curves) and with the addition of the second-order derivatives (solid red curves), respectively.
-The lower panel in each plot gives the ratio of the two uncertainties.
-IV Pseudo global analysis with new data
-The eigenvectors in a global analysis are determined from a particular set of data inputs.
-Therefore, both the first-order and second-order uncertainty estimations depend on the eigenvectors.
-To further test the improved Hessian method when the eigenvectors change, we introduce new experimental observables into the global fit, which are the new proton structure parameters $P_{u}$ and $P_{d}$ factorized from the forward-backward asymmetry of the Drell-Yan process [8], and the boost asymmetry $A_{\text{boost}}$ in the diboson process [9].
-These asymmetry variables were recently proposed at hadron colliders which can provide unique information on the proton structure.
-The $P_{u}$ and $P_{d}$ parameters factorize the proton structure information in $pp(u\bar{u})\rightarrow Z/\gamma^{*}\rightarrow\ell^{+}\ell^{-}$ and $pp(d\bar{d})\rightarrow Z/\gamma^{*}\rightarrow\ell^{+}\ell^{-}$ events.
-The two light quark initial states can be decoupled because the electroweak couplings of $u$-$\bar{u}$-$Z$ and $d$-$\bar{d}$-$Z$ are different.
-As a result, $P_{u}$ and $P_{d}$ separately provide the $u$ and $d$ quark information, which are mixed and indistinguishable in the measurement of the total cross section $\sigma^{pp}_{Z}$.
-$A_{\text{boost}}$ is the asymmetry defined in the diboson production [9].
-For example, in the $pp(u\bar{d})\rightarrow W^{+}\gamma\rightarrow\ell^{+}\nu+\gamma$ process, the $W^{+}$ boson predominantly couples to the $\bar{d}$ quark while the photon couples to the $u$ quark.
-Consequently, the boost asymmetry $A_{\rm boost}^{W^{+}\gamma}$ has a positive large value, namely $\gamma$ is more boosted than the $W^{+}$ boson, because the initial state parton lunminosity $u(x_{1}){\bar{d}}(x_{2})$ is larger than $u(x_{2}){\bar{d}}(x_{1})$ for $x_{1}>x_{2}$, etc.
-Therefore, the relative asymmetry between the initial states $q_{i}(u_{1})\bar{d}(x_{2})$ and $u(x_{2})\bar{d}(x_{1})$ ($x_{1}>x_{2}$), which arises from the nature of the valence and sea quark differences, can be measured by comparing the energy or rapidity of the two bosons.
-According to Ref. [8] and Ref. [9], the new pseudo observables can be constructed as:
-$P_{u}(x_{1},x_{2})$ $=$ $\left[u_{V}(x_{1})+\bar{u}(x_{1})\right]\bar{u}(x_{2})-\left[u_{V}(x_{2})+\bar{u}(x_{2})\right]\bar{u}(x_{1})$
-$P_{d}(x_{1},x_{2})$ $=$ $\left[d_{V}(x_{1})+\bar{d}(x_{1})\right]\bar{d}(x_{2})-\left[d_{V}(x_{2})+\bar{d}(x_{2})\right]\bar{d}(x_{1})$
-$A^{W^{+}\gamma}_{\text{boost}}(x_{1},x_{2})$ $=$ $\left[u_{V}(x_{1})+\bar{u}(x_{1})\right]\bar{d}(x_{2})-\left[u_{V}(x_{2})+\bar{u}(x_{2})\right]\bar{d}(x_{1})$
-$A^{W^{-}\gamma}_{\text{boost}}(x_{1},x_{2})$ $=$ $\left[d_{V}(x_{1})+\bar{d}(x_{1})\right]\bar{u}(x_{2})-\left[d_{V}(x_{2})+\bar{d}(x_{2})\right]\bar{u}(x_{1})$
-To gain information on the $x$-dependence of the PDFs, the above pseudo observables are calculated with the $|Y|$ values the same as the previous settings.
-The relative uncertainty of each data point is also assigned as $3\%$.
-We repeat the calculations of the Hessian matrix now, using both the cross section observables and the new asymmetry observables; so we acquire the eigenvectors, eigenvalues, and the new orthogonal parameters $\{z_{i}\}$ accordingly.
-The new eigenvalues are shown in Fig.
+Following the method described in section IV, the uncertainty of any physical observable $\mathcal{O}$ can be estimated according to Eq.
+(19).
+To numerically study the difference between the original Hessian method and the improved Hessian method, we compare the relative uncertainties of $xf_{q}(x)$ for $u_{V}$, $d_{V}$, $\bar{u}$, $\bar{d}$ and $s$, which are shown in Fig.
 5.
-With more data constraints introduced, the eigenvalues are in general larger.
-In Fig.
-6, $\Delta\chi^{2}=\chi^{2}(z_{i}=1)-\chi^{2}_{\text{min}}$ is shown both in the previous pseudo global fit and the new fit.
-As discussed, $\Delta\chi^{2}=1$ indicates a good quadratic approximation, so that the corresponding eigenvector is well constrained by data.
-Therefore, more eigenvectors have $\Delta\chi^{2}=1$ or close to 1 in the new global fit, as expected.
-Figure 5: 24 eigenvalues in the pseudo global analysis with new asymmetry observables included.
-Figure 6: $\Delta\chi^{2}=\chi^{2}(z_{i}=1)-\chi^{2}_{\text{min}}$ in the pseudo global fit with new asymmetry observables included, compared with the previous global fit with only cross section observables.
-The relative uncertainties on $xf_{q}(x)$ for $u_{V}$, $d_{V}$, $\bar{u}$, $\bar{d}$ and $s$ are shown in Fig.
-7.
-As the constraints from the input data are strengthened, the estimated uncertainties are reduced with respect to the previous global fit in Fig.
-4.
-In principle, when the uncertainties of individual eigenvectors decrease, the second-order uncertainties in Eq.
-(1) become less important with respect to the first-order ones.
-In Appendix B, we give another pseudo global fit in which the relative uncertainty of each data point is reduced to $0.1\%$.
-With such a high precision, there is indeed no difference between the first-order and second-order calculation, as expected.
-In the present section, although the new asymmetry observables provide stronger constraints, their precisions are not high enough to reduce the second-order terms to a negligible level.
-The ratios between the first-order and second-order uncertainties in Fig.
-7 can become even larger than the previous global fit in Fig.
-4, due to the complexity of the multiple data fit.
-This reflects the fact that the uncertainty estimation in a PDF global fit depends highly on the data inputs.
-Therefore, the second-order calculation is not only a necessary procedure to get a better uncertainty estimation, but also a sufficient way to evaluate the power of the data constraints in a global fit.
-When the difference between the first-order and second-order uncertainties is small, it indicates a strong data constraint so that the output PDF is less dependent on various assumptions and the choice of non-perturbative functions.
-Figure 7: Relative uncertainty on $xf_{u_{V}}(x)$, $xf_{d_{V}}(x)$, $xf_{\bar{u}}(x)$, $xf_{\bar{d}}(x)$ and $xf_{s}(x)$.
-The new asymmetry observables $P_{u}$, $P_{d}$, and $A^{W^{\pm}\gamma}_{\text{boost}}$ are included in the global analysis.
-V Conclusion
-In this work, we propose an improved Hessian method in the PDF global analysis, which estimates the uncertainty with the inclusion of the complete set of second-order derivatives.
-A numerical test has been designed, in which a set of PDFs with multiple parameters are fitted to pseudo experimental observables.
-The test shows that when the precision of the input data used in a global analysis is limited, the improved Hessian method would give different uncertainties compared to the original method.
-Since the difference in the uncertainty can be reduced as the data constraints become stronger, one can tell whether a PDF global analysis is experimentally well constrained and less model-dependent, by comparing the uncertainties of the two methods.
+Figure 5: Relative uncertainties on $xf_{u_{V}}(x)$, $xf_{d_{V}}(x)$, $xf_{\bar{u}}(x)$, $xf_{\bar{d}}(x)$ and $xf_{s}(x)$.
+Comparison includes the contribution of the first-order derivatives which is estimated from the original Hessian method (blue curves), the uncertainty of $\delta\mathcal{O}_{E}$ which estimates the non-linearity of $T_{\alpha}(a_{i})$ (green curves), the uncertainty of the second-order derivatives additional to the first-order contributions (orange curves), and the total uncertainty from the improved Hessian method which is the quadratic sum of the previous three (red curves).
+The lower panel in each plot gives the ratio of the total uncertainty from the improved Hessian method to the original Hessian method.
+Fig.
+5 shows that the non-linear uncertainty in this pseudo global analysis is much larger than the original uncertainty, especially in the large and small $x$ region where no data can provide direct constraint.
+Even though the pseudo data inputs in this work are generated using the non-perturbative functions exactly the same as that used in the pseudo theory calculations, the uncertainty of the choice of non-perturbative functions can still be reflected in the non-constrained area.
+The DIS-liked observables are dominated by the valence quarks of $u_{V}$ and $d_{V}$ around the peak region of their distributions at $x\sim 0.1$, where the contributions of the sea quarks are relatively low.
+For the sea quarks however, there is no data predominated by them.
+Therefore, the valence pseudo quark PDFs are constrained more directly than the sea quark PDFs in this study.
+Consequently, the non-linear uncertainties for $u_{V}$ and $d_{V}$ are generally smaller than that for the sea quarks, as shown in Fig.
+5.
+Moreover, the second-order uncertainty on the parton densities is not as significant as the non-linear uncertainty.
+If a physical observable contains information from multiple partons, the second-order uncertainties could be larger.
+In practice however, it could be fine not to compute the second-order uncertainty with thousands of second-order error sets for simplicity, especially when the non-linear uncertainty is already quoted.
+In this work, the total uncertainties in the pseudo global analysis is several times larger than the original ones.
+According to the study from the CTEQ-TEA PDFs, the original uncertainties have to be enlarged corresponding to $\Delta\chi^{2}=37$, at the $68\%$ confidence level, to cover the difference of using various non-perturbative functions, which also leads to a total uncertainty several times larger.
+The MSHT20 PDF gives a similar conclusion.
+In their studies, the PDF uncertainty is defined by a dynamic tolerance which effectively corresponds to $\Delta\chi^{2}\sim 10$, giving the final uncertainty about 3 times larger than the original one.
+As a conclusion, the non-linear effect in the Hessian method is essential according to today’s experimental precisions.
 VI Acknowledgement
@@
 PHY-2310291.
-Appendix A Appendix A: PDF-induced uncertainty on physical observables
-In this section, we give detailed calculation of Eq. (14). Rewrite Eq. (1) as:
+Appendix A Appendix: PDF-induced uncertainty on physical observables
+In this section, we give the detailed calculation of Eq. (17). Rewrite Eq. (6) as:
 $\mathcal{O}(z_{i})$ $=$ $\mathcal{O}(z^{0}_{i})+\sum_{i}A_{i}+\frac{1}{2}\sum_{i\neq j}B_{ij}$
@@
 Furthermore, including only the second-order diagonal terms generally underestimates the uncertainty.
-Appendix B Appendix B: A pseudo global fit with high precision data
-A pseudo global fit with high precision data is performed in this section.
-9 data variables are used, including $\sigma^{\text{DIS}}_{\gamma}$, $\sigma^{\text{DIS}}_{Z}$, $\sigma^{pp}_{Z}$, $\sigma^{pp}_{W^{+}}$, $\sigma^{pp}_{W^{-}}$, $P_{u}$, $P_{d}$, $A^{W^{+}\gamma}_{\text{boost}}$ and $A^{W^{-}\gamma}_{\text{boost}}$.
-For each data point corresponding to a particular $x$ value, the relative uncertainty is set to $0.1\%$.
-With sych high precision data used in the global fit, the uncertainty of the second-order derivative calculation is consistent with the first-order calculation.
-The result is shown in Fig.
-8.
-Figure 8: Relative uncertainty on $xf_{u_{V}}(x)$, $xf_{d_{V}}(x)$, $xf_{\bar{u}}(x)$, $xf_{\bar{d}}(x)$ and $xf_{s}(x)$.
-The new asymmetry observables $P_{u}$, $P_{d}$, and $A^{W^{\pm}\gamma}_{\text{boost}}$ are included in the global analysis.
-Relative uncertainty of each data point is improved to $0.1\%$.